Problem 13-5 Bonus compensation; algebra [L013-3 Sometimes compensation packages include bonuses designed to provide performance incentives to employees. The difficulty a bonus can cause accountants is not an accounting problem, but a math problem. The complication is that the bonus formula sometimes specifies that the calculation of the bonus is based in part on the bonus itself. This occurs anytime the bonus is a percentage of income because expenses are components of income, and the bonus is an expense. Regalia Fashions has an incentive compensation plan through which a division manager receives a bonus equal to 8% of the divisions net income. Division income in 2018 before the bonus and income tax was $110,000 The tax rate is 30%. Required 2. Calculate the amount of the bonus. 3. Prepare the adjusting entry to record the bonus compensation. 4. Bonus arrangements take many forms. Suppose the bonus specifies that the bonus is 8% of the divisions income before tak but after the bonus itself. Calculate the amount of the bonus. Answer is not complete. Complete this question by entering your answers in the tabs below Required 2 Required 3 Required 4 Calculate the amount of the bonus. (Do not round intermediate calculations and round your answers to the nearest whole dollar amount.) $ 972 3
Answer #1

(2)

Income before bonus and tax = $110,000

Let income after bonus and tax (net income) be = $Y

Bonus = 8% of net income

Hence, bonus = Y x 8%

= 0.08Y

Income after bonus = 110,000 - 0.08Y

Income tax rate = 30%

Hence, income tax = (110,000 - 0.08Y) x 30%

= 33,000 - 0.024Y

Net income = Income before bonus and tax - Bonus - Income tax

Y = 110,000 - 0.08Y - 33,000 + 0.024Y

1.056Y = 77,000

Y = 77,000/1.056

= $72,917

Net income = $72,917

Bonus = 0.08Y

= 0.08 x 72,917

= $5,833

Hence, bonus amount is $5,833

(3)

Journal

Date Account title Debit Credit
Bonus expense 5,833
Cash 5,833

(4)

Income before bonus and tax = $110,000

Bonus = 8% of income after bonus

Hence, bonus = 110,000 x 8/108

= $8,148

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.